OpenStudy (jenniferjuice):

HELLL00000 I NEEEDDDD HELPPPPPP?????? Solve 3^x−5 = 9. x = −3 x = 2 x = 7 x = 8

OpenStudy (ikram002p):

Hello im here :P

OpenStudy (jenniferjuice):

:P please help

OpenStudy (anonymous):

im here toooooooooooooo

OpenStudy (jenniferjuice):

@Kayla_Is_Smart aawsome :D

OpenStudy (ikram002p):

wait ur Qn like this \(\Huge 3^{x-5}=9\)

OpenStudy (ikram002p):

hint 3^2=9

OpenStudy (anonymous):

lol @jenniferjuice

OpenStudy (ikram002p):

\(\Huge 3^{x-5}=3^2\) x-5= 2 solve for x :)

OpenStudy (jenniferjuice):

9=9?

OpenStudy (ikram002p):

sigh x-5=2 x=2+5 x=7

OpenStudy (jenniferjuice):

thank u

OpenStudy (ikram002p):

np :)